# Hallmarks of Optimization Trajectories in Neural Networks and LLMs: The Lengths, Bends, and Dead Ends

### v1

@@ 3 A Tale of Hyperparameters
4 A Ride with Momentum, Decay, & More
In order to obtain a more refined understanding of the above observation, we would like to inspect in detail the angular and norm-based measures of the trajectory. Let’s start by taking a look at how aligned are the net or the aggregate updates [^4] We qualify this by ‘net’ or ‘aggregate’ as we are working on somewhat coarser granularity (i.e., $1$ epoch) than actual updates or steps. But our current granularity is still rich enough to allow for the presented trends to persist even if go 2×2$\times$2 × to 5×5$\times$5 × more coarser. in the presence of these hyperparameters and otherwise.
@@ 4.1 Momentum and the Angle between Updates
In particular, in Figure 4(a), we plot the angle between consecutive epochs, i.e, $\angle(\theta_{t+1}-\theta_{t},\theta_{t}-\theta_{t-1})$. Interestingly, we find that this angle becomes obtuse a short while into the training process, and as a matter of fact, this angle is larger when momentum is turned on versus when off. Moreover, rather visibly, this increase in angle is larger when weight decay is also enabled, suggesting that weight decay and momentum are closely intertwined. Taken as such, this observation would point as to how the MDS increases when these hyperparameters are switched off. But, we would still like to gain a somewhat better understanding of this mechanism, even though doing so in precise details might be out of reach. Hence, we turn to the simplest and oft-employed model of a quadratic problem, and see where do we arrive.
@@
Given a quadratic problem with $\ell_{2}$ regularization of strength $\alpha>0$, namely,
$\min_{{\boldsymbol{\theta}}\in{\mathbb{R}^{d}}}\frac{1}{2}{\boldsymbol{\theta}}^{\top}{\bf M}\,{\boldsymbol{\theta}}+\frac{1}{2}\alpha\|{\boldsymbol{\theta}}\|^{2}\,,$
with ${\bf M}\in{\mathbb{R}^{d\times d}}$ symmetric with eigenvalues $\lambda_{1}\geq\cdots\lambda_{d}$, the angle between successive steps $\Delta_{t}={\boldsymbol{\theta}}_{t}-{\boldsymbol{\theta}}_{t-1},\Delta_{t+1}={\boldsymbol{\theta}}_{t+1}-{\boldsymbol{\theta}}_{t}$, when using gradient descent with a one-step momentum ($\mu>0$) and learning rates $\eta_{t},\eta_{t+1}$, can be upper and lower bounded as follows:
$\langle\Delta_{t},\Delta_{t+1}\rangle\leq\eta_{t}\eta_{t+1}(1-\mu\eta_{t}-\eta_{t}\alpha-\eta_{t}\lambda_{d})(\lambda_{d}+\alpha)^{2}\|{\boldsymbol{\theta}}_{t-1}\|^{2}$
$\langle\Delta_{t},\Delta_{t+1}\rangle\geq\eta_{t}\eta_{t+1}(1-\mu\eta_{t}-\eta_{t}\alpha-\eta_{t}\lambda_{1})(\lambda_{1}+\alpha)^{2}\|{\boldsymbol{\theta}}_{t-1}\|^{2}$
Proof.
See Appendix A. ∎
The above proof inherently considers the solution at ${{\boldsymbol{\theta}}^{\star}}=\mathbf{0}$, but if that is not the case, we can substitute for it in the objective and our derived upper and lower bounds would scale in the squared distance to the solution, i.e. $\|{\boldsymbol{\theta}}_{t-1}-{{\boldsymbol{\theta}}^{\star}}\|^{2}$. Besides, in the above proof, we consider a one-step momentum, which inherently means resetting the momentum after every $2$ steps. This is done for convenience, as our main purpose is to anyways gain insights into the phenomenon and not provide its ultimate proof.
Turning to the bounds themselves, notice that if the learning rate $\eta_{t}\geq 1/(\lambda_{1}+\mu+\alpha)$, the lower bound will turn negative and will be multiplied by a factor of $(\lambda_{1}+\alpha)^{2}\|{\boldsymbol{\theta}}_{t-1}\|^{2}$. On the other hand, while the first term of the upper bound might still be positive, importantly, it is scaled by a factor of $(\lambda_{d}+\alpha)^{2}\approx\alpha^{2}$ for matrices ${\bf M}$ which are close to degenerate.
@@
In our context, the Hessian of the loss with respect to the parameters will play the role of the matrix ${\bf M}$, since we can assume a second-order Taylor series will hold across the two steps. But it is also known through prior empirical work that the Hessian is significantly degenerate (Sagun et al., 2017), which has also been proven rigorously for deep linear fully-connected and convolutional networks (Singh et al., 2021). Furthermore, this requirement on the learning rate $\eta_{t}$ is actually looser than the adaptivity of the largest eigenvalue of the Hessian to the learning rate $\lambda\approx\frac{2}{\eta}$, as shown in the recent work on Edge of Stability (EoS) (Cohen et al., 2022).
@@
Owing to these facts, we will have that $\lambda_{d}({\bf M})\approx 0$, and which further implies that the upper bound on the inner-product between the updates will be approximately zero, and the lower-bound will be large in absolute value and negative. Therefore, this explains how the angles between consecutive epochs can turn out to be obtuse. More broadly, the obtuse angle indeed implies that there are oscillations, especially along the direction of the largest Hessian eigenvector. Further, from Lemma 1, we see that the magnitude of the inner-product of the updates scales in proportion to $\eta_{t+1}$. Hence, a way to dampen the oscillations [^5] These oscillations need not necessarily translate into gross instabilities at the level of the loss, since as can be seen in Figure 4(g), the update norms progressively shrink in each of the three learning rate phases. is to decrease the learning rate, and as can be seen in Figure 4(a), the learning rate decay at epochs $30$ and $60$ is followed right after with the angles turning from obtuse to acute. Lastly, here in the constraint on the learning rate (the additive terms $\alpha$ and $\mu$), we can also see how momentum and weight decay go hand-in-hand, each accentuating the effect of the other.
Figure 5: Trajectory Maps of ResNet50 for varying models across different amounts of Weight Decay, with the corresponding MDS scores being $\omega=0.731,0.679,0.844,0.882,0.885$.
@@ 4.2 More than just Weight Decay
Now that we have gained a richer understanding of momentum and its interaction with weight decay, let us turn to weight decay alone and understand its sole effect on the directional nature of the trajectories. We already noticed the increase in mean directional similarity (MDS) when weight decay is disabled for ResNet50, trained with SGD on ImageNet. In fact, we find a similar effect if switch to an adaptive optimizer, like AdamW (Loshchilov & Hutter, 2017), the trajectory maps for which can be found in Figure 5. Here, we used regularization constants from $\lambda=0$ until the first value where we witness a decrease in test performance, which in this case was $\lambda=1$. Specifically, we analyze the weight decay coefficients in $\lambda\in\{1,0.1,0.01,0.001,0\}$. The corresponding MDS come out to be, $\omega=0.731,0.679,0.844,0.882,0.885$.
We notice that, as before, increasing weight decay leads to a heightened directional exploration, or lower MDS. Except for the case of $\lambda=1$, which seems to be a slight anomaly. If we take a look at its corresponding trajectory map in Figure 5, we find that there is a much darker square grid that onsets around epoch $30$. But also here we find increased cosine similarities in a larger band around the diagonal, relative to the case of $\lambda=0.1$. As it turns out, with such a high weight decay coefficient $\lambda=1$, this particular network underfits, achieving a train/test top-1 accuracy as $54.63\%,50.52\%$ respectively. This fact can also be confirmed by the plots of distance from initialization and the apex angle at initialization, which are located in Appendix C.3. The performance for the rest of the networks improves, more or less, as expected with weight decay, and in particular, achieve accuracies of $75.45\%,73.38\%,71.03\%,71.41\%$. Furthermore, in such a case of underfitting, we notice that computing the relative MDS from initialization corrects for the underfitting itself, resulting in the values of $\omega_{0}$ as $0.985,0.807,0.862,0.897,0.900$ for $\lambda=1,0.1,0.01,0.001,0$ respectively.
Having reaffirmed our results extensively for this setting, let us understand how directional exploration is encouraged by weight decay. The Figure 7 gives a physics-based intuition into this aspect. In particular, we can think of the loss gradient pulling the network parameters rightwards, while the force exerted by weight decay tries to pull the network downwards. The relative strengths of these two ‘forces’ have been represented by the lengths of the two vector arrows. We notice that as the weight decay strength is increased, from the left subfigure to the right, the angle traced at the origin (O) also increases. This quite simply explains how weight decay can contribute towards directional exploration.
@@
Figure 7: Directional Exploration due to Weight Decay. The pink downward vector represents the pull towards the origin (O) due to an energy corresponding to the norm squared of the parameters, while the leftward vector in purple represents the force due to the objective energy (classification loss).
4.3 Directional Effects in Other Key Settings
In addition to the momentum and weight decay, there are other crucial hyperparameters, such as learning rate and batch size, whose directional effects warrant a mention. We carry out additional experiments in these settings, and from where, the key findings are that the learning rate, as it would be easy to guess, indeed encourages directional exploration leading to low MDS scores. But, somewhat more interestingly, we find that increasing the batch size also helps further exploration and thereby decreases the MDS scores. The trajectory maps can be found in the Figure 6 While we encourage the curious reader to have a look at the Appendix C.11, we find that with increased batch size, the angle between the updates as well as the angle between the update and the current location become increasingly obtuse, and thus making room for a wider directional exploration. In contrast, for smaller batch sizes these angles are closer to $90^{\circ}$. We hypothesize that a similar mutual interaction, as observed with weight decay and momentum, also occurs with batch size is considered. A detailed analysis, however, remains outside the current scope.
Lastly, we also experimented with Sharpness-Aware Minimization (Foret et al., 2021) (SAM), where we found that a higher value of the SAM regularization coefficient leads to a slightly increased directional similarity, which could potentially be related to SAM directing optimisation to flatter basins wherein the individual points are more directionally alike and have higher cosine similarities. The detailed results can be found in the Appendix C.5.
Other Settings and Datasets.
5 Hauling Trajectory Hallmarks for LLMs
Having covered rather extensively, the crucial set of hyperparameters as well as optimizers, the one question that would still be on the minds of readers: how does it behave with scale? With all the massive improvements wrought in by Large Language Models (LLMs) and their impressive scaling behaviour, it would be relevant to uncover the structure of trajectories across models of different sizes. Several questions arise: (a) Do we see similar patterns in trajectories of networks trained on language modelling tasks? (b) What is their trend over increasing model sizes? (c) Do we expect increasing parameter scale to provide new directions for learning or do we expect scale to regularize the occurrence of complex trajectories?
@@ 5 Hauling Trajectory Hallmarks for LLMs
Thanks to Pythia’s (Biderman et al., 2023) publically released model checkpoints over training, for GPT-NeoX (Black et al., 2022) models — ranging in sizes from $14$ Million (M) to $12$ Billion (B) — we can provide answers into the above questions. Given that processing all of the available checkpoints would require several terabytes of cache, we select every fourth checkpoint, resulting in $39$ checkpoints that we analyze for models of sizes: $14$M, $70$M, $160$M, $410$M, $1.4$B, $2.8$B, $6.9$B, $12$B. The results for these experiments can be found in Figure 8, with the corresponding MDS being: $\omega=0.650,0.672,0.678,0.726,0.759,0.786,0.818,0.815$.
@@
First, we note that there is a tiny square grid in the upper-left corner, which delineates precisely the learning rate warmup phase. Next, the subsequent larger grid starts out, for $14$M, with distinct subgrids but then over increasing model scale takes a funnel-like shape around $410$M, before becoming rather homogeneous by $12$B parameters. Even the horizontal and vertical slivers corresponding to warmup and the rest of the epochs start to assume a higher cosine similarity with scale. Overall, increasing scale lends an intense dark hue to the trajectory maps, which seems to offer a concrete hypothesis the inductive bias of scale might be related to regularizing the trajectories.
@@
We show that this surprising finding about the progressive increase in cosine similarity with scale has a relatively simple explanation, at least in the case of the large-width limit of deep networks. The gist of our argument is that in the large width limit, any parameter updates that lead to stable feature updates must necessarily yield updated parameters that are identically aligned with their initialisation. This is a well-known fact for lazy learning regimes like the Neural Tangent Kernel (Jacot et al., 2020) or standard parameterisations, where no feature learning occurs. What may be surprising is that this is necessarily true for feature learning regimes like $\mu$(Yang et al., 2022). The precise details can be found in Appendix B.
@@
Figure 9: Layerwise Trajectory Maps for Q,K,V weights across depth fot GPT-NeoX trained on Pile for the $14$M model.
6 Related Work and Discussion

### v2

@@ 4.1 Momentum and the Angle between Updates
In particular, in Figure 4(a), we plot the angle between consecutive epochs, i.e, $\angle(\theta_{t+1}-\theta_{t},\theta_{t}-\theta_{t-1})$. Interestingly, we find that this angle becomes obtuse a short while into the training process, and further, this angle is larger when momentum is turned on versus when off. Moreover, rather visibly, this increase in angle is larger when weight decay is also enabled, suggesting that weight decay and momentum are closely intertwined. Taken as such, this observation would point as to how the MDS increases when these hyperparameters are switched off. To gain a somewhat better understanding of this mechanism, we turn to the simplest and oft-employed model of a quadratic problem.
@@
Given a quadratic problem with $\ell_{2}$ regularization of strength $\alpha>0$, namely, $\min_{{\boldsymbol{\theta}}\in{\mathbb{R}^{d}}}\frac{1}{2}{\boldsymbol{\theta}}^{\top}{\bf M}\,{\boldsymbol{\theta}}+\frac{1}{2}\alpha\|{\boldsymbol{\theta}}\|^{2}\,,$ with ${\bf M}\in{\mathbb{R}^{d\times d}}$ symmetric with eigenvalues $\lambda_{1}\geq\cdots\geq\lambda_{d}$, the angle between successive steps $\Delta_{t}={\boldsymbol{\theta}}_{t}-{\boldsymbol{\theta}}_{t-1},\Delta_{t+1}={\boldsymbol{\theta}}_{t+1}-{\boldsymbol{\theta}}_{t}$, when using gradient descent with a one-step momentum ($\mu>0$) and learning rates $\eta_{t},\eta_{t+1}$, can be upper and lower bounded as follows:
$\langle\Delta_{t},\Delta_{t+1}\rangle\leq\eta_{t}\eta_{t+1}(1-\eta_{t}(\mu+\alpha+\lambda_{d}))(\lambda_{d}+\alpha)^{2}\|{\boldsymbol{\theta}}_{t-1}\|^{2}$
$\langle\Delta_{t},\Delta_{t+1}\rangle\geq\eta_{t}\eta_{t+1}(1-\eta_{t}(\mu+\alpha+\lambda_{1}))(\lambda_{1}+\alpha)^{2}\|{\boldsymbol{\theta}}_{t-1}\|^{2}$
The proof, in Appendix A, inherently considers the solution at ${{\boldsymbol{\theta}}^{\star}}=\mathbf{0}$, but if that is not the case, we can substitute it in the objective and our derived bounds would scale in the squared distance to the solution, i.e. $\|{\boldsymbol{\theta}}_{t-1}-{{\boldsymbol{\theta}}^{\star}}\|^{2}$. Besides, in the above proof, we consider a one-step momentum, which inherently means resetting the momentum after every $2$ steps. This is done for convenience, as our main purpose is to anyways gain insights into the phenomenon and not provide its ultimate proof.
Turning to the bounds themselves, notice that if the learning rate $\eta_{t}\geq 1/(\lambda_{1}+\mu+\alpha)$, the lower bound will turn negative and will be multiplied by a factor of $(\lambda_{1}+\alpha)^{2}\|{\boldsymbol{\theta}}_{t-1}\|^{2}$. On the other hand, although the first term of the upper bound might still be positive, importantly, it is scaled by a factor of $(\lambda_{d}+\alpha)^{2}\approx\alpha^{2}$ for matrices ${\bf M}$ which are close to degenerate ($\lambda_{d}\rightarrow 0$).
@@
In our context, the Hessian of the loss with respect to the parameters will play the role of the matrix ${\bf M}$, since we can assume a second-order Taylor series will hold across the two steps. But it is also known through prior empirical work that the Hessian is significantly degenerate [Sagun et al., 2017], which has also been proven rigorously for deep linear fully-connected and convolutional networks [Singh et al., 2021]. Furthermore, this requirement on the learning rate $\eta_{t}$ is actually looser than the adaptivity of the largest eigenvalue of the Hessian to the learning rate $\lambda\approx\frac{2}{\eta}$, as shown in the recent work on Edge of Stability (EoS) [Cohen et al., 2022].
@@
Owing to these facts, we will have that $\lambda_{d}({\bf M})\approx 0$, and which further implies that the upper bound on the inner-product between the updates will be approximately zero, and the lower-bound will be large in absolute value but negative. Therefore, this explains how the angles between consecutive epochs can be obtuse. More broadly, the obtuse angle indeed implies that there are oscillations, especially along the direction of the largest Hessian eigenvector. Further, from Lemma 1, we see that the magnitude of the inner-product of the updates scales in proportion to $\eta_{t+1}$. Hence, a way to dampen the oscillations [^7] These oscillations need not necessarily translate into gross instabilities at the level of the loss, since as can be seen in Figure 4(g), the update norms progressively shrink in each of the three learning rate phases. is to decrease the learning rate, and as can be seen in Figure 4(a), the learning rate decay at epochs $30$ and $60$ is followed right after with the angles turning from obtuse to acute. Lastly, here in the constraint on the learning rate (the additive terms $\alpha$ and $\mu$), we can also see momentum and weight decay go hand-in-hand, each accentuating the effect of the other.
@@ 4.2 More than just Weight Decay
Figure 5: Relative Trajectory Maps (wrt. initialization) of ResNet50 models for different weight decay.
@@
Figure 6: Directional exploration effect of Weight Decay. The downward vector represents the pull towards the origin (O) due to weight decay, while the rightward vector the force due to the loss .
Now that we have gained a richer understanding of momentum and its interaction with weight decay, let us turn to weight decay alone and understand its directional effect. We have already noticed the increase in mean directional similarity (MDS) when weight decay is disabled for ResNet50 trained with SGD on ImageNet. In fact, we find a similar effect with an adaptive optimizer, like AdamW [Loshchilov and Hutter, 2017] — the trajectory maps for which are shown in Figure 5. Here, we used regularization constants from $\lambda=0$ until the first value where we witness a decrease in test performance, which in this case was $\lambda=1$. Specifically, we analyze the weight decay coefficients in $\lambda\in\{1,0.1,0.01,0.001,0\}$. The corresponding MDS come out to be, $\omega=0.731,0.679,0.844,0.882,0.885$. We notice that, as before, increasing weight decay leads to a heightened directional exploration, or lower MDS; except $\lambda=1$ being the seeming anomaly.
But we find that this can be remedied simply by looking at the relative trajectory maps (Figure 5), and computing the relative MDS, i.e., $\omega_{0}$ is $0.985,0.807,0.862,0.897,0.900$ for $\lambda=1,0.1,0.01,0.001,0$ respectively. This occurs since such a high weight decay $\lambda=1$, causes this particular network to underfit (train/test top-1 accuracy are $54.63\%,50.52\%$). The performance for the rest of the networks improves, more or less, as expected with weight decay, and in particular, achieve accuracies of $75.45\%,73.38\%,71.03\%,71.41\%$.
Having reaffirmed our results extensively about the directional exploration due to weight decay, we can understand it through a simple physics-based intuition, as shown in the Figure 6. In particular, we can think of the loss gradient pulling the network parameters rightwards, while the force exerted by weight decay tries to pull the network downwards. The relative strengths of these two ‘forces’ have been represented by the lengths of the two vector arrows. We notice that as the weight decay strength is increased, from the left subfigure to the right, the angle traced at the origin (O) also increases. This explains how weight decay can contribute towards directional exploration.
5 Hauling Trajectory Hallmarks for LLMs and their model internals
Figure 7: Trajectory Maps of Pythia GPT-NeoX models across two orders of scales trained on Pile. The corresponding MDS $\omega=0.650,0.678,0.759,0.815$.
With all recent interest in Large Language Models (LLMs) and their scaling, a natural question is whether the structure of trajectories exhibited before is shared in the case of language modelling tasks and across models of different sizes. Besides, while we expect increasing parameter count to provide new directions for learning, does scale make the optimization trajectories complex or does it instead regularize them?
@@ 5 Hauling Trajectory Hallmarks for LLMs and their model internals
Thanks to Pythia’s [Biderman et al., 2023] publically released model checkpoints over training, for GPT-NeoX [Black et al., 2022] models — ranging in sizes from $14$ Million (M) to $12$ Billion (B) — we can provide answers to the above questions. Given that processing all the available checkpoints would require several terabytes of cache, we select every fourth checkpoint, resulting in $39$ checkpoints that we analyze for models of sizes: $14$M, $70$M, $160$M, $410$M, $1.4$B, $2.8$B, $6.9$B, $12$B. The results for a shortlist of these experiments can be found in Figure 7 (for more, see Figure 69).
@@
First, we note that there is a tiny square grid in the upper-left corner, which delineates precisely the learning rate warmup phase. Next, the subsequent larger grid starts out, for $14$M, with distinct subgrids but then with increasing model scale takes a funnel-like shape around $1.4$B, before becoming rather homogeneous by $12$B parameters. Even the horizontal and vertical slivers corresponding to warmup and the rest of the epochs start to assume a higher cosine similarity with scale. Overall, increasing scale lends an intense dark hue to the trajectory maps, which seems to suggest that the inductive bias of scale might be related to regularizing the trajectories.
@@
We prove, in Appendix B, that this surprising finding about the progressive increase in cosine similarity with scale has a relatively simple explanation, at least in the case of the large-width limit of deep networks. The gist of our argument is that in the large width limit, any parameter updates that lead to stable feature updates must necessarily yield updated parameters that are identically aligned with their initialisation. This is a well-known fact for lazy learning regimes like the Neural Tangent Kernel [Jacot et al., 2020] or standard parameterisations, where no feature learning occurs. What may be surprising is that this is necessarily true for feature learning regimes like $\mu$[Yang et al., 2022].
Layerwise Q,K,V dynamics homogenize over depth at $\sim 1$B scale and over.
Another interesting aspect of the structure of trajectory maps for the $14$M and $160$M cases is their relative heterogeneity (distinct subgrids), as compared to the billion parameter models. This heterogeneity cannot be explained away through the settings of a learning rate scheduler (which is just a cosine scheduler), so we inspect the layerwise trajectory maps (i.e., by building it over parameters of each layer separately). In particular, we find an intriguing heterogeneous structure for these models that is most starkly present in the query-key-value (Q,K,V) parameters of the attention layers, especially the middle layers, as shown in Figure 8.
@@
Figure 8: Layerwise Trajectory Maps for Q,K,V weights across depth fot GPT-NeoX trained on the Pile for the $160$M model. This suggests that the Q,K,V parameters of the middle layers seem to be converging last, and differ from earlier and later layers in this regard.
(a) $14$ M
(b) $160$ M
(c) $1.4$ B
(d) $12$ B
Figure 9: Trajectory maps of Q,K,V layers become homogenized over increasing scale.
This indicates that the Q,K,V parameter dynamics converge at different timescales, with the middle layers converging the last directionally in contrast to earlier and later layers. Moreover, we find that scale has the striking effect of homogenizing the Q,K,V dynamics, as shown in Figure 9, with the layerwise trajectory maps essentially resembling the network-wide trajectory map structure in Figure 7.
6 Putting Directional Redundancy to the Test
Our trajectory map analyses reveal significant directional redundancy in the optimization trajectories of neural networks, which is especially prominent later into training. This raises the question of whether this redundancy can be leveraged for creating efficient hybrid optimization schemes or if it conceals minute but crucial directional changes. Therefore, to test the true nature of directional redundancy, we consider optimizing only some layerwise scalar parameters after a little while into training. Further, instead of attaching extra scalar parameters per neuron or channel and tuning them, we will repurpose the scalar weight and bias parameters present in the ubiquitous (batch/layer) normalization layers in modern network architectures. We are inspired by the prior work of Frankle et al. [2021] who demonstrated the remarking expressivity of training just the scalar parameters in batch-normalization (BN) layers in the entire network.
CIFAR10 experiments.
We experiment with ResNet20 on CIFAR10 using SGD over 160 epochs, freezing non-BN layers at different points and training only the $1,376$ scalar parameters in the BN layers. Figure 10 presents the results alongside the trajectory map of the full network training.
Figure 10: Directional Redundancy put to test by training only the BN scalar parameters. The trajectory map of the original, full network, training is shown on the left and, in the bar chart to the right, the bars denoting the performance achieved by training BN parameters from a particular epoch are horizontally aligned with the corresponding trajectory map rows. ‘Baseline’ denotes the network accuracy just before switching to optimizing the BN parameters only.
When right from the initialization, just the BN layers are trained, as claimed in Frankle et al. [2021], the resulting network achieves a test accuracy of $54.8\%$, which although quite interesting falls nevertheless considerably short of the $91.2\%$ test accuracy obtained by training the entire network. In a way, from the horizontal strip of trajectory map around 00, which pales as we move from the left to right, we see that the full network parameters at the end of training deviate significantly in their directionality as opposed to the parameters at initialization (the cosine similarity being $\sim 0.3$). However soon after, from around $40$ epochs, where the trajectory map starts developing a dark hue (and the cosine similarity to the final parameters climbs to about $\sim 0.9$), training BN parameters alone brings us to within $2\%$ of the full-network accuracy; and from around $80$ epoch gets to within $1\%$, and completely matches thereafter. Notably, this feat is remarkably achieved by training only $0.5\%$ of the overall parameters, and this fares even better than training all the parameters in the bulkier last layer as shown in Figure 77.
ImageNet experiments.
Likewise, in our experiments on ResNet50 trained on ImageNet as in Figure 2, we found that although training BN layers right from initialization gets us to an accuracy of just $\sim 6.4\%$ (a paltry amount, given the full-network’s accuracy of $\sim 76\%$), but training it from epoch $30$ gets to within $\sim 10\%$ of the full-network’s accuracy and from epoch $60$ to within about $\sim 2\%$ of it, i.e, to $\sim 73.2\%$ top-1 test accuracy. This is again a striking result, since when training the BN parameters of ResNet50, only $0.18\%$ of the parameters are being used and $99.82\%$ of the parameters are kept frozen.
It should be mentioned that in the above BN training procedure, although the forward and backward propagation runtime costs end up being similar to training the entire network, this still leads to savings in the GPU memory consumption as the optimization buffers (like the momentum buffer or those used for preconditioning in adaptive methods like Adam) now need to be of significantly smaller size (typically $<1\%$, i.e., $99\%$ savings) to accommodate only the BN parameters and could, in turn, make larger batch sizes feasible. More broadly, we foresee that this idea can be readily adapted into a hybrid optimization scheme, where regular training can be interleaved with memory-light training of normalization layer parameters, and where the latter is done on cheaper GPU instances in the cloud, hence contributing to non-trivial cost savings. All in all, these findings pointedly show that the observed directional redundancy from the trajectory maps truly manifests during training and can be potentially utilized to reap practical gains.
7 Related Work